# Negative impact of heavy-tailed uncertainty and error distributions on the reliability of calibration statistics for machine learning regression tasks

### v3

@@ Abstract
Average calibration of the prediction uncertainties of machine learning regression tasks can be tested in two ways: one is to estimate the calibration error (CE) as the difference between the mean absolute error (MSE) and the mean variance (MV) or mean squared uncertainty; the alternative is to compare the mean squared z-scores (ZMS) or scaled errors to 1. The problem is that both approaches might lead to different conclusions, as illustrated in this study for an ensemble of datasets from the recent machine learning uncertainty quantification (ML-UQ) literature. It is shown that the estimation of MV, MSE and their confidence intervals can become unreliable for heavy-tailed uncertainty and error distributions, which seems to be a common issue for ML-UQ datasets. By contrast, the ZMS statistic is less sensitive and offers the most reliable approach in this context. Unfortunately, the same problem affects also conditional calibrations statistics, such as the popular ENCE, and very likely post-hoc calibration methods based on similar statistics. As not much can be done to relieve this issue, except for a change of paradigm to intervals- or distribution-based UQ metrics, robust tailedness metrics are proposed to detect the potentially problematic datasets.
@@ I Introduction
The next section defines the calibration statistics and the validation approach. Sect. III introduces the ML-UQ datasets used to illustrate the validation results, with a focus on the shape of their uncertainty and error distributions. Numerical experiments, based on synthetic datasets mimicking the real datasets, are performed in Sect. IV, in order establish the impact of the tailedness of the uncertainty and error distributions on the reliability of the calibration statistics. Sect. V illustrates this reliability issue on the reference ML-UQ datasets for RCE and ZMS, and estimates its impact on conditional calibration statistics. The main conclusions are presented in Sect. VI.
@@ II.2 $ZMS$ and related statistics
Another approach to calibration based on Eq. 1 uses scaled errors or z-scores
@@ II.3 Validation
which considers explicitly the possible asymmetry of $I_{BS}$ around $\vartheta_{est}$. The compatibility of the statistic with its reference value can then be tested by
@@ III.1 $u_{E}^{2}$ distributions
where $\alpha>0$ and $\beta>0$ are the shape and scale parameters, respectively, expressed as combinations of a number of degrees of freedom $\nu$ and a scale factor $\sigma$. This equation is more conveniently expressed as
@@
This distributions is often used in Bayesian inference as a conjugate prior for the variance in a normal likelihood.
Figure 1: Fit of the squared errors (histogram) by an Inverse-Gamma $\Gamma^{-1}(\nu,\nu)$ distribution (blue line).
Fit of squared uncertainty distributions by the IG model is done by maximum goodness-of-fit estimation using the Kolmogorov-Smirnov distance (Delignette2015, ). The results are shown in Fig. 1, and the shape parameters reported in Table 1.
Overall, the fits are rather good. The distributions for Sets 1, 2 and 6 are the worst ones, mostly due to the presence of a bimodality in these datasets (a tiny one for Set 2). The $\nu$ values for these sets are very low: for Set 2 $\nu<1$ corresponds to an IG distributions with undefined mean value, whereas for Sets 1 and 6 values of $\nu<2$ indicate an undefined variance. These values should not be over-interpreted. Nevertheless, better fitted distributions present also very small shape parameters, such as Sets 4 and 7, meaning that such values might not be extravagant.
@@ III.2 $E^{2}$ distributions
Figure 2: Fit of the squared errors (histogram) by a Fisher-Snedecor $F(1,\nu)$ distribution (blue line). The red curves represent the distributions for NIG model compatible with Fig. 1.
The squared error datasets have been fitted by a scaled $F$ distribution, and the results are reported in Fig. 2 and Table 1. The fits are very good for all sets, except for Set 6, where a minor mode at small errors is not accounted for by the $F$ distribution. Here again, the shape parameters cover a wide range, from 0.8 to 10.
It has to be noted that, although the fits are good, the shape parameter for $E^{2}$ is never close to be twice the shape parameter for $u_{E}^{2}$ as suggested by the NIG model. The densities expected from the NIG model as plotted in Fig. 2 (red lines) for comparison with the actual fits and confirm the discrepancy. Therefore, either the generative distribution $D$ is never close to a normal for the studied datasets, or the datasets are not properly calibrated.
@@ III.3 $Z^{2}$ distributions
The shape parameters for the fit of $Z^{2}$ distributions by the same procedure as for $E^{2}$ are reported in Table 1. The fits present the same features as $E^{2}$ fits (not shown), albeit with typically larger shape parameters, ranging from 0.9 to 20, which is expected from the generative model. Two exceptions are Sets 5 and 8, for which the shape parameter is slightly smaller for $Z^{2}$ than for $E^{2}$, which might not be a significant difference.
@@ IV.1 Alternative tailedness metrics
The shape of the uncertainty and error distributions can rather well be characterized by a shape parameter $\nu$, but one has to take into account that not all datasets are correctly fitted by the model distributions. Besides, for the $IG$ and $t$ distributions, shape metrics such as skewness and kurtosis are not defined for $\nu\leq 3$ or 4, respectively. One therefore needs distribution-agnostic metrics able to characterize the shape of these distributions.
@@
Fig. 3 shows the correspondence between $\nu$, $\beta_{GM}$ and $\kappa_{CS}$ for samples of the IG distribution [$X^{2}\sim\Gamma^{-1}(\nu/2,\nu/2)$] and F-distribution [$X^{2}\sim F(1,\nu)$]. There is a monotonous correspondence between $\beta_{GM}$ , $\kappa_{CS}$ and the $\nu$ parameter of the sampled distributions showing that the skewness and kurtosis statistics can be used to replace unambiguously the shape parameter of the model distributions.
@@
Note that the coefficient of variation $c_{v}$ is often used in ML-UQ studies to quantify the dispersion of uncertainties(Scalia2020, ), which is also somehow related to the tailedness of their distribution. However, it is not robust and cannot replace the proposed statistics. In the following, $\beta_{GM}$ and $\kappa_{CS}$ are used (and compared) to characterize the tailedness of the distributions.
@@ IV.2 Sensitivity of calibration statistics to heavy-tailed uncertainty and error distributions
To characterize the sensitivity of the RCE and ZMS statistics to the shape of the uncertainty and error distributions in a validation context, a comparison of the rates of failure of validation tests for both statistics is performed on synthetic datasets. $N=10^{3}$ synthetic calibrated datasets of size $M=5000$ are generated using two scenarios designed to cover the range of distributions observed in the previous section:
  • NIG - a normal generative distribution [$D=N(0,1)$] and an Inverse Gamma distribution for $u_{E}^{2}$ [$u_{E}^{2}\sim\Gamma^{-1}(\nu_{IG}/2,\nu_{IG}/2)$], with $\nu_{IG}$ varying between 2 and 10. Values of $\nu_{IG}$ smaller than 2 have been avoided as they might result in numerical problems.
  • TIG - a scaled Student’s-$t$ generative distribution
    $D=t_{s}(\nu_{D})\coloneqq t(\nu_{D})\sqrt{\frac{\nu_{D}-2}{\nu_{D}}}$ (20)
    with $\nu_{D}$ varying between 2.1 and 20, and a fixed Inverse Gamma distribution for $u_{E}^{2}$ [$u_{E}^{2}\sim\Gamma^{-1}(3,3)$]. This scenario aims to reproduce the error distributions with small shape parameters $\nu$ that are not accessible directly by the NIG model.
For each sample, the calibration is tested by $|\zeta|\leq 1$ and a probability of validity is estimated as
@@
The values of $p_{val}$ and their 95% confidence intervals obtained by a binomial model(Pernot2022a, ), are plotted in Fig. 4 for the RCE and ZMS statistics as a function of $\nu_{IG}$ or $\nu_{D}$. The upper axis provides the average $\beta_{GM}$ and $\kappa_{CS}$ statistics for the generated $u_{E}^{2}$ and $E^{2}$ samples.
@@
Figure 4: Validation probability of the ZMS and RCE statistics for calibrated datasets generated by two scenarios: (left) NIG with $\nu_{IG}$ as parameter of the IG distribution; (right) TIG with $\nu_{D}$ as parameter of the generative $D=t_{s}$ distribution. The corresponding values of $\beta_{GM}$ and $\kappa_{CS}$ are reported on the upper axis.
@@
The study of the NIG model shows that the ZMS is not sensitive to $\nu_{IG}$, even for extreme uncertainty distributions, and that it provides intervals that consistently validate 95 % of the calibrated synthetic datasets. For the RCE, the validation error is strongly sensitive to $\nu_{IG}$ for values below 4, reaching more than 20 % for $\nu_{IG}=2$. When considering the TIG model, one sees that both statistics are similarly sensitive to the shape of the generative distribution, and that the validation intervals begin to be unreliable for $\nu_{D}<6$. The failure rate reaches over 30 % for $\nu_{D}=3$. In this case, when $\nu_{D}$ diminishes, the dispersion of the generated $<E^{2}>$ or $<Z^{2}>$ values increases, which is not in itself a problem, but the bootstrapped CIs are unable to capture correctly this effect by being too narrow, i.e. too frequently excluding the true mean value of the distribution.
@@
For further reference one defines safety thresholds for the skewness and kurtosis metrics. Based on the data in Fig. 4, the limit values above which one can suspect the reliability of the calibration statistics are given in Table 2. Note that these values are indicative and result from the distribution shapes observed for the studied datasets. The values for other distribution shapes might slightly differ.
Variable $\beta_{GM}$ $\kappa_{CS}$ Impacted stats
$u_{E}^{2}$ 0.6 3.0 RCE
$E^{2}$, $Z^{2}$ 0.8 5.0 RCE, ZMS
@@
Table 2: Limits for the skewness and kurtosis metrics above which the reliability of the RCE or ZMS metrics should be questioned.
IV.3 Summary
@@
This section was designed to illustrate how the calibration statistics and their validation diagnostic are sensitive to the shape of the uncertainty and error distributions. The main problem is the sensitivity of the mean square (MS) statistic $<X^{2}>$, which is central to the studied RCE and ZMS, to the presence of outliers or heavy tails in the distribution of $X$. This is a well known issue, which justifies often the replacement of the root mean squared error RMSE by the more robust mean unsigned error (MUE) in performance analysis(Pernot2018, ; Pernot2021, ). Unfortunately the use of the MS in average calibration statistics derives directly from the probabilistic model linking errors to uncertainties (Eq. 1), and one has to deal with its limitations, unless one is ready to change of uncertainty paradigm, using for instance prediction intervals or distributions.
@@
The robust skewness ($\beta_{GM}$) and kurtosis ($\kappa_{CS}$) estimators were proposed and compared here as tailedness metrics, and safety limits for both metrics have been derived to help in screening out datasets with potentially unreliable RCE/ZMS values. It has been shown that heavy-tailed uncertainty distributions affect mostly the RCE, while both RCE and ZMS are affected by heavy-tailed error distributions. The ZMS benefits also from the fact that the distribution of $Z^{2}$ has often lighter tails than the distribution of $E^{2}$. The next section illustrates these features on the nine example ML-UQ datasets.
V Application to the ML-UQ datasets
@@
The validation approach presented above is applied to the datasets presented in Sec. III. First, the datasets are characterized by their skewness/kurtosis values to validate the analysis of shape parameters reported above. Then, the comparison of calibration diagnostics for the RCE and ZMS is shown and analyzed. Finally, the impact on binned calibration statistics is evaluated.
V.1 Skewness and kurtosis analysis
@@
The first step is to assess the tailedness of the uncertainty, error and z-score distributions to reveal potentially problematic datasets. The corresponding $\beta_{GM}$ and $\kappa_{CS}$ values are reported in Table 3.
Set # $\beta_{GM}(u_{E}^{2})$ $\kappa_{CS}(u_{E}^{2})$ $\beta_{GM}(E^{2})$ $\kappa_{CS}(E^{2})$ $\beta_{GM}(Z^{2})$ $\kappa_{CS}(Z^{2})$
1 0.40 -0.20 0.82 5.06 0.73 2.32
2 0.72 4.10 0.94 19.68 0.83 6.37
3 0.66 3.19 0.74 2.19 0.69 1.48
4 0.74 5.67 0.82 4.52 0.69 2.07
5 0.19 1.84 0.78 4.32 0.79 4.07
6 0.50 1.46 0.96 22.70 0.95 23.97
7 0.93 3.91 0.98 9.84 0.78 3.97
8 0.30 0.41 0.79 4.77 0.78 4.69
9 0.30 0.48 0.77 5.06 0.75 4.48
@@
Table 3: Robust skewness and kurtosis values for $u_{E}^{2}$, $E^{2}$ and $Z^{2}$. Values above the proposed safety limits (Table 2) are in bold type.
The skewness and kurtosis metrics provide essentially identical diagnostics about potentially problematic datasets, the exception being Set 9 which has an error distribution slightly below the skewness limit and slightly above the kurtosis limit. For the remaining sets, one might concentrate on skewness.
@@
The $\beta_{GM}$ values reported in Table 3 indicate that some distributions present heavy tails. For uncertainties, $\beta_{GM}(u_{E}^{2})$ exceed 0.6 for Sets 2, 3, 4, and 7. This list conforms to the shape analysis of Table 1, except for Set 6, which had a very small $\nu$ parameter but has a moderate skewness. For the errors, the largest values occur for Sets 1, 2, 4, 6 and 7, a list where the shape analysis of Table 1 would have added Set 5. For the z-scores, only two sets are screened out, Sets 2 and 6, while Sets 5 and 6 have small $\nu$ values. The correspondence between the previous shape analysis and the skewness/kurtosis screening is not perfect, which reflects the imperfect quality of the distributions fits provided by the IG and F models. According to the present analysis, Sets 1, 2, 3, 4, 6 and 7 are potentially problematic, Set 1 being very close to the limits for both skewness and kurtosis.
@@
A graphical summary of the skewness and kurtosis analysis is provided in Fig. 5. Skewness is represented in Fig. 5(left), where $\beta_{GM}(E^{2})$ is plotted against $\beta_{GM}(u_{E}^{2})$ for the nine datasets (red triangles) and for the NIG model ($2\leq\nu\leq 20$; black dots). One sees directly that only Sets 5, 8 and 9 are in the safe area for both statistics, although they lie close to the $E^{2}$ skewness limit. All the other sets have a skewness value that might be problematic for at least one of $u_{E}^{2}$ or $E^{2}$. It is also noteworthy that most $\beta_{GM}(E^{2})$ values (except for Sets 3, 4 and 7) exceed what should be expected from the NIG model, confirming the shape analysis of Sect. III.
@@
Figure 5: Skewness (left) and kurtosis (right) analysis of the application datasets. The black dots figure the NIG model for $2\leq\nu\leq 20$. The colored areas signal the values that might be associated with problems.
@@
The blue squares depict the values for $\beta_{GM}(Z^{2})$: except for Sets 2 and 6, they all lie below the safety limit (0.8), and for Set 2 the value is much closer to the limit than $\beta_{GM}(E^{2})$. This indicates that the ZMS is much less likely to be affected by estimation problems than the RCE, except maybe for Sets 2 and 6.
The kurtosis analysis provides essentially the same results [Fig. 5(right)], except for points very close to a limit.

### v4

@@ Abstract
Average calibration of the (variance-based) prediction uncertainties of machine learning regression tasks can be tested in two ways: one is to estimate the calibration error (CE) as the difference between the mean absolute error (MSE) and the mean variance (MV); the alternative is to compare the mean squared z-scores (ZMS) to 1. The problem is that both approaches might lead to different conclusions, as illustrated in this study for an ensemble of datasets from the recent machine learning uncertainty quantification (ML-UQ) literature. It is shown that the estimation of MV, MSE and their confidence intervals becomes unreliable for heavy-tailed uncertainty and error distributions, which seems to be a frequent feature of ML-UQ datasets. By contrast, the ZMS statistic is less sensitive and offers the most reliable approach in this context. Unfortunately, the same problem is expected to affect also conditional calibrations statistics, such as the popular ENCE, and very likely post-hoc calibration methods based on similar statistics. Several solutions to circumvent the outlined problems are proposed.
@@ I Introduction
The next section defines the calibration statistics and the validation approach. Sect. III introduces the ML-UQ datasets used to illustrate the validation results, with a focus on the shape of their uncertainty and error distributions. Numerical experiments, based on synthetic datasets mimicking the real datasets, are performed in Sect. IV, in order establish the impact of the tailedness of the uncertainty and error distributions on the reliability of the calibration statistics. Sect. V illustrates this reliability issue on the reference ML-UQ datasets for RCE and ZMS. The main conclusions are presented in Sect. VI, along with a proposition of solutions.
@@ II.2 $ZMS$ and related statistics
Another approach to calibration based on Eq. 1 uses z-scores
@@ II.3 Validation
which considers explicitly the asymmetry of $I_{BS}$ around $\vartheta_{est}$. The compatibility of the statistic with its reference value can then be tested by
@@ III.1 $u_{E}^{2}$ distributions
where $\alpha>0$ and $\beta>0$ are the shape and scale parameters, respectively, expressed here as combinations of a number of degrees of freedom $\nu$ and a scale factor $\sigma$. This equation is more conveniently expressed as
@@
This distributions is often used in Bayesian inference as a prior for variance parameters.
Figure 1: Fit of the squared uncertainties (histogram) by an Inverse-Gamma $\Gamma^{-1}(\nu,\nu)$ distribution (blue line).
Fit of squared uncertainty distributions by the IG model is done by maximum goodness-of-fit estimation using the Kolmogorov-Smirnov distance (Delignette2015, ). The results are shown in Fig. 1, and the shape parameters are reported in Table 1.
Overall, the fits are of contrasted quality. The distributions for Sets 1, 2 and 6 are the worst ones, mostly due to the presence of a bimodality (a tiny one for Set 2). The $\nu$ values for these sets are very low: for Set 2 $\nu<1$ corresponds to an IG distributions with undefined mean value, whereas for Sets 1 and 6 values of $\nu<2$ indicate an undefined variance. These values should not be over-interpreted. Nevertheless, better fitted distributions present also very small shape parameters ($\nu<2$), such as Sets 4 and 7, meaning that such values might not be unreasonable.
@@ III.2 $E^{2}$ distributions
Figure 2: Fit of the squared errors (histogram) by a Fisher-Snedecor $F(1,\nu)$ distribution (blue line). The red curves represent the distributions for NIG models compatible with Fig. 1.
The squared error datasets have been fitted by a scaled $F$ distribution, and the results are reported in Fig. 2 and Table 1. The fits are very good for all sets, except for Set 6, where a minor mode at small errors cannot be accounted for by the $F$ distribution. Here again, the shape parameters cover a wide range, from 0.8 to 10, with very values typical of heavy tails.
It has to be noted that the shape parameter for $E^{2}$ is never close to being twice the shape parameter for $u_{E}^{2}$ as expected from the NIG model (Eq. 18). The densities expected from the NIG model are plotted in Fig. 2 (red lines) for comparison with the actual fits and confirm the discrepancy. Therefore, either the generative distribution $D$ is never close to $N(0,1)$ for the studied datasets, or the datasets are not properly calibrated.
@@ III.3 $Z^{2}$ distributions
The shape parameters for the fit of $Z^{2}$ distributions by the same procedure as for $E^{2}$ are reported in Table 1. The fits (not shown) present the same features as $E^{2}$ fits, albeit with typically larger shape parameters, ranging from 0.9 to 20, which is expected from the generative model. Two exceptions are Sets 5 and 8, for which the shape parameter is slightly smaller for $Z^{2}$ than for $E^{2}$, a small difference that might not be significant.
@@ IV.1 Alternative tailedness metrics
As shown in the previous section, the shape of the uncertainty and error distributions can be rather well characterized by a shape parameter $\nu$, but one has to take into account that not all datasets are correctly fitted by the model distributions. Besides, for the $IG$ and $t$ distributions, shape metrics such as skewness and kurtosis are not defined for $\nu\leq 3$ or 4, respectively. One therefore needs distribution-agnostic metrics able to characterize the shape of these distributions.
@@
Fig. 3 shows the correspondence between $\nu$, $\beta_{GM}$ and $\kappa_{CS}$ for samples of the IG distribution [$X^{2}\sim\Gamma^{-1}(\nu/2,\nu/2)$] and F-distribution [$X^{2}\sim F(1,\nu)$]. There is a monotonous one-to-one correspondence between $\beta_{GM}$ , $\kappa_{CS}$ and the $\nu$ parameter of the sampled distributions showing that the skewness and kurtosis statistics can be used to replace unambiguously the shape parameter of the model distributions.
@@
Note that the coefficient of variation $c_{v}$ is often used in ML-UQ studies to quantify the dispersion of uncertainties(Scalia2020, ), which is also related to the tailedness of their distribution. However, it is not robust and cannot replace the proposed statistics.
Considering that $\beta_{GM}$ and $\kappa_{CS}$ are practically exchangeable for the kind of distributions considered here, only $\beta_{GM}$ is used in the following to characterize the tailedness of the distributions. It provides a conveniently bounded metric.
@@ IV.2 Sensitivity of calibration statistics to heavy-tailed uncertainty and error distributions
To characterize the sensitivity of the RCE and ZMS statistics to the shape of the uncertainty and error distributions, they are estimated for synthetic datasets covering the range of tail statistics observed in the previous sections. $N=10^{3}$ synthetic calibrated datasets of size $M=5000$ are generated using an extension of the NIG model, the TIG model, defined by $u_{E}^{2}\sim\Gamma^{-1}(\nu_{IG}/2,\nu_{IG}/2)$ and $D=t_{s}(\nu_{D})$) where
$t_{s}(\nu_{D})\coloneqq t(\nu_{D})\times\sqrt{\frac{\nu_{D}-2}{\nu_{D}}}$ (20)
[…]
The RCE deviates to positive values when $\beta_{GM}(E^{2})>0.85$, and the deviation is increasing with skewness. As the Monte Carlo error bars are not large enough to cover the reference value (0), one can conclude that the RCE sis significantly biased in this area. Note that reformulating the RCE without square roots can help to reduce this bias (see Appendix A). In comparison, the 1-ZMS statistic is better behaved for most of the datasets, and seems to have problems only with the extreme $\nu_{D}=3$ case, which corresponds to $\beta_{GM}(Z^{2})>0.8$.
IV.3 Sensitivity of validation statistics to heavy-tailed uncertainty and error distributions
The previous experiment revealed how the RCE and ZMS are sensitive to heavy tails. It occurs that the estimation of confidence intervals for these statistics by bootstrapping has also difficulties in similar conditions, which is put in evidence by the following simulations.
One considers here two scenarios designed to cover the range of distributions observed above, but also to limit the computational cost of the experiments:
  • NIG: $u_{E}^{2}\sim\Gamma^{-1}(\nu_{IG}/2,\nu_{IG}/2)$ and $D=N(0,1)$, with $\nu_{IG}$ varying between 2 and 10.
  • TIG: $u_{E}^{2}\sim\Gamma^{-1}(3,3)$ and $D=t_{s}(\nu_{D})$ with $\nu_{D}$ varying between 2.1 and 20.
For each sample ($M=5000$ with $N=10^{3}$ Monte Carlo repeats), the calibration is tested by $|\zeta|\leq 1$ and a probability of validity is estimated as
@@ IV.3 Sensitivity of validation statistics to heavy-tailed uncertainty and error distributions
The values of $p_{val}$ and their 95% confidence intervals obtained by a binomial model(Pernot2022a, ), are plotted in Fig. 5 for the RCE and ZMS statistics as a function of $\nu_{IG}$ or $\nu_{D}$. The upper axis provides the average $\beta_{GM}$ statistics for the generated $u_{E}^{2}$ and $E^{2}$ samples.
@@
Figure 5: Validation probability of the ZMS and RCE statistics for calibrated datasets generated by two scenarios: (left) NIG with $\nu_{IG}$ as parameter of the IG distribution; (right) TIG with $\nu_{D}$ as parameter of the generative $D=t_{s}$ distribution. The corresponding average values of $\beta_{GM}$ are reported on the upper axis.
The study of the NIG model shows that the ZMS is not sensitive to $\nu_{IG}$, even for extreme uncertainty distributions, and that it provides intervals that consistently validate 95 % of the calibrated synthetic datasets. For the RCE, the validation probability is strongly sensitive to $\nu_{IG}$ for values below 4, reaching less than 80 % for $\nu_{IG}=2$.
@@
When considering the TIG model, one sees that both statistics are similarly sensitive to the shape of the generative distribution, and that the validation intervals begin to be unreliable for $\nu_{D}<6$ ($\beta_{GM}\geq 0.8$). The validation probability falls to 0.65 for $\nu_{D}=2.1$. In this case, when $\nu_{D}$ diminishes, the dispersion of the generated $<E^{2}>$ or $<Z^{2}>$ values increases, which is not in itself a problem, but the bootstrapped CIs are unable to capture correctly this effect by being too narrow, i.e. too frequently excluding the true mean value of the distribution.
@@
For further reference one defines safety thresholds for $\beta_{GM}$. Based on the data in Figs. 4-5, the limit values above which one can suspect the reliability of the calibration statistics are given in Table 2. Note that these values are indicative and result from the distribution shapes observed for the studied datasets. The values for other distribution shapes might differ slightly.
Variable $\beta_{GM}$ Impacted statistics
$u_{E}^{2}$ 0.6 RCE
$E^{2}$, $Z^{2}$ 0.8 RCE, ZMS
@@
Table 2: Limits for $\beta_{GM}$ above which the reliability of the RCE or ZMS calibration statistics should be questioned.
IV.4 Summary
This section was designed to illustrate how the calibration statistics and their validation diagnostic are sensitive to the shape of the uncertainty and error distributions. Two problems were identified.
@@ IV.4 Summary
The first problem is the sensitivity of the mean square (MS) statistic $<X^{2}>$, which is central to the studied RCE and ZMS, to the presence of outliers or heavy tails in the distribution of $X$. This is a well known issue, which justifies often the replacement of the root mean squared error RMSE by the more robust mean unsigned error (MUE) in performance analysis(Pernot2018, ; Pernot2021, ). Unfortunately the use of the MS in average calibration statistics derives directly from the probabilistic model linking errors to uncertainties (Eq. 1), and one has to deal with its limitations, unless one is ready to change of uncertainty paradigm, using for instance prediction intervals or distributions.
The second problem is the inability of bootstrapping to provide reliable confidence intervals for $<X^{2}>$ for heavy-tailed distributions.
@@
The robust skewness ($\beta_{GM}$) and kurtosis ($\kappa_{CS}$) estimators were proposed and compared as tailedness metrics. Because of their strong correlation, only $\beta_{GM}$ was retained and safety limits have been derived to help in screening out datasets with potentially unreliable RCE/ZMS values. It has been shown that heavy-tailed uncertainty distributions affect mostly the RCE, while both RCE and ZMS are affected by heavy-tailed error distributions. The ZMS benefits also from the fact that the distribution of $Z^{2}$ has often lighter tails than the distribution of $E^{2}$. The next section illustrates these features on the nine example ML-UQ datasets of Sect. III.
V Application to real ML-UQ datasets
@@
The validation approach presented above is applied to the datasets presented in Sec. III. First, the datasets are characterized by their skewness to validate the analysis of shape parameters reported above. Then, the comparison of calibration diagnostics for the RCE and ZMS is analyzed according to the skewness values.
V.1 Skewness analysis
@@
The first step is to assess the tailedness of the uncertainty, error and z-score distributions to reveal potentially problematic datasets. The corresponding $\beta_{GM}$ values are reported in Table 3.
Set # $\beta_{GM}(u_{E}^{2})$ $\beta_{GM}(E^{2})$ $\beta_{GM}(Z^{2})$
1 0.40 0.82 0.73
2 0.72 0.94 0.83
3 0.66 0.74 0.69
4 0.74 0.82 0.69
5 0.19 0.78 0.79
6 0.50 0.96 0.95
7 0.93 0.98 0.78
8 0.30 0.79 0.78
9 0.30 0.77 0.75
@@
Table 3: Robust skewness values for $u_{E}^{2}$, $E^{2}$ and $Z^{2}$. Values above the proposed safety limits (Table 2) are in bold type.
@@
The skewness values indicate that some distributions present heavy tails. For uncertainties, $\beta_{GM}(u_{E}^{2})$ exceed 0.6 for Sets 2, 3, 4, and 7. This list conforms to the shape analysis of Table 1, except for Set 6, which had a very small $\nu$ parameter but has a moderate skewness. For the errors, the largest values occur for Sets 1, 2, 4, 6 and 7, a list where the shape analysis of Table 1 would have added Set 5. For the z-scores, only two sets are screened out, Sets 2 and 6, while Sets 5 and 6 have small $\nu$ values. The correspondence between the previous shape analysis and the skewness screening is not perfect, which reflects the imperfect quality of the distributions fits provided by the IG and F models. According to the present analysis, Sets 1, 2, 3, 4, 6 and 7 are potentially problematic, Set 1 being very close to the limit.
@@
A graphical summary of the skewness analysis is provided in Fig. 6, where $\beta_{GM}(E^{2})$ is plotted against $\beta_{GM}(u_{E}^{2})$ for the nine datasets (red triangles) and for the NIG model ($2\leq\nu\leq 20$; black dots). One sees directly that only Sets 5, 8 and 9 are in the safe area for both statistics, although they lie close to the $E^{2}$ skewness limit. All the other sets have a skewness value that might be problematic for at least one of $u_{E}^{2}$ or $E^{2}$. It is also noteworthy that most $\beta_{GM}(E^{2})$ values (except for Sets 3, 4 and 7) exceed what should be expected from the NIG model, confirming the shape analysis of Sect. III.
@@
Figure 6: Skewness analysis of the application datasets. The black dots figure the NIG model for $2\leq\nu\leq 20$. The colored areas signal the values that might be associated with problems.
@@
The blue squares depict the values for $\beta_{GM}(Z^{2})$: except for Sets 2 and 6, they all lie below the safety limit (0.8), and for Set 2 the value is much closer to the limit than $\beta_{GM}(E^{2})$. This indicates that the ZMS is much less likely to be affected by estimation problems than the RCE, but might still be challenged by data such as Sets 2 and 6.